# thanks: Corresponding author

### v1

@@ IV.3 EA of Non-valence correlation-bound Anions
The additional electron in the non-valence correlation-bound ((NVCB)) anions are weakly bound in a diffuse orbital by means of electron correlation. Due to their sensitivity, NVCB anions serve as a challenging case for natural orbital based approximate wave-function method. There are several molecules extensively studied in the literature that contain NVCB anionic state eg. TCNE, [vooraTheoreticalApproachesTreating2017] polycyclic aromatic hydrocarbons, [vooraNonvalenceCorrelationBoundAnion2015] perfluorobenzene, [vooraNonvalenceCorrelationBoundAnion2014] fullerenes [vooraNonvalenceCorrelationBoundAnion2014a] etc. In this work, we have studied the NVCB anionic state of perfluorobenzene ((C6F6)).
Perfluorobenzene has a diffuse nonvalence correlation-bound anionic state in its planar equilibrium geometry. This state evolves barrierlessly into the valence-bound anion as the molecule buckles $(C_{2v})$ and may contribute to the high electron mobility in liquid C6F6[vooraNonvalenceCorrelationBoundAnion2014] The geometry of C6F6 is optimized in ORCA [neeseSoftwareUpdateORCA2022] using RI-MP2/aug-cc-pVDZ. The optimized geometry is provided in the Supplementary Material.
The EA value corresponding to the NVCB anion is calculated using the aug-cc-pVTZ basis set with 7s7p diffuse functions on a ghost atom at the center of the molecule. The aug-cc-pVTZ/C auxiliary basis has been used. The results are presented in TABLE 4. It can be seen that the EA-ADC(2) method significantly overestimates the EA value compared to the EOM-CCSD and EOM-MP2 results. The EA-ADC(3) method, on the other hand, underestimates the EA value. The sm-EA-ADC[(2)+x(3)] method gives excellent agreement with the available EOM results. The convergence of EA values is much slower for the NVCB state as compared to the valence states in both ADC(3) and sm-EA-ADC[(2)+x(3)] as the additional electron is distributed over the entire molecules. It can be seen that at the threshold of $10^{-4}$, the sm-EA-ADC[(2)+x(3)] shows an error of 0.05 eV compared to the canonical result. The error reduces to 0.03 eV on going to the $10^{-5}$. The error does not completely converge even on going to the threshold of $10^{-6}$ and shows a negligible error of 0.02 eV.
@@ IV.4 Computational Efficiency
The computational efficiency of the SS-FNO-EA-ADC(3) method was further evaluated by calculating the electron affinity of a large molecular system, Zn-protoporphyrin, which has been previously investigated for electron attachment. [duttaPairNaturalOrbital2016] The molecular geometry was taken from the Supplementary Material of Ref. duttaPairNaturalOrbital2016. The calculation was performed on a dedicated workstation equipped with an Intel(R) Xeon(R) E5-2667 v4 @ 3.20 GHz processor and 512 GB of RAM. The aug-cc-pVTZ basis set together with the aug-cc-pVTZ/C auxiliary basis was employed for the Zn atom, while the aug-cc-pVDZ and aug-cc-pVDZ/C basis sets were used for all other atoms. Zn-protoporphyrin consists of 75 atoms and 326 electrons, with a total of 1184 virtual orbitals. Application of the SS-FNO truncation threshold of $10^{-4}$ reduced the virtual orbital space to 807, and the number of natural auxiliary functions decreased from 4053 to 3440 with the NAF threshold of $10^{-2}$. The total computation time for the SS-FNO-EA-ADC(3) calculation was 1 day, 10 hours, and 13 minutes, with the SCF step requiring only 18 minutes. The canonical ADC(2) calculation for a single root required 2 hours and 8 minutes, of which 34 minutes were spent on intermediate generation. In the truncated basis, the ADC(3) calculation took 1 day, 7 hours, and 20 minutes, with 17 hours and 49 minutes required for intermediate construction. The ADC(2) calculation took 20 minutes in the SS-FNO basis. The SS-FNO-EA-ADC(3) method gives an electron affinity of 1.041 eV for the lowest electron-attached state. The corresponding Dyson orbital for Zn-protoporphyrin is depicted in FIG. 10.
@@ V Conclusion
We have developed, implemented, and benchmarked a reduced-cost ADC(3) approach for the computation of electron affinities, based on the state-specific frozen natural orbital (SS-FNO) framework within the non-Dyson intermediate state representation (ISR) formalism. Unlike the conventional MP2-based FNO scheme, the state-specific FNO approach selects an optimal subset of virtual natural orbitals tailored for each electron-attached state, thereby achieving a substantial reduction in computational cost for ADC(3) calculations. To further enhance efficiency, the density-fitting approximation has been employed to eliminate the need for storing four-index electron repulsion integrals, and an additional truncation of the auxiliary basis via natural auxiliary functions (NAFs) provides further computational savings.
@@ Acknowledgements.
This work has been supported by the SERB-India under the CRG and MATRICS schemes, awarded to AKD. AKD acknowledges the research fellowship funded by the EU NextGenerationEU through the Recovery and Resilience Plan for Slovakia under project No. 09I03-03-V04-00117.

### v2

@@ IV.3 EA of Non-valence correlation-bound Anions
The additional electron in the non-valence correlation-bound ((NVCB)) state is weakly bound in a diffuse orbital through electron correlation. Due to their sensitivity, NVCB anions provide a challenging test case for natural-orbital-based approximate wave-function methods. Several molecules containing NVCB anionic states have been extensively studied in the literature, for example TCNE, [78] polycyclic aromatic hydrocarbons, [77] perfluorobenzene, [75] fullerenes [76] etc. IIn this work, we study the NVCB anionic state of perfluorobenzene (C6F6).
Perfluorobenzene possesses a diffuse non-valence correlation-bound anionic state at its planar equilibrium geometry. This state evolves barrierlessly into a valence-bound anion as the molecule buckles to a $C_{2v}$ geometry and may contribute to the high electron mobility observed in liquid C6F6[75] The geometry of C6F6 was optimized using ORCA [48] using RI-MP2/aug-cc-pVDZ. at the RI-MP2/aug-cc-pVDZ level of theory.
The electron affinity corresponding to the NVCB anion was calculated using the aug-cc-pVTZ basis set with additional $7s7p$ diffuse functions placed on a ghost atom at the center of the molecule. The results are presented in Table 4. The EA-ADC(2) method significantly overestimates the electron affinity compared with the EOM-CCSD and EOM-MP2 results. In contrast, the EA-ADC(3) method underestimates the electron affinity. The convergence of EA values is much slower for the NVCB state as compared to the valence states in both ADC(3) and sm-EA-ADC[(2)+x(3)] as the additional electron is distributed over the entire molecule in the former. At a truncation threshold of $10^{-4}$, the sm-EA-ADC[(2)+x(3)] method shows an error of approximately 0.05 eV relative to the canonical result. The error decreases to approximately 0.03 eV when the threshold is tightened to $10^{-5}$. Even at a threshold of $10^{-6}$, the error does not fully converge but remains small at approximately 0.02 eV.
@@ IV.4 Computational Efficiency
The computational efficiency of the SS-FNO-EA-ADC(3) method was further evaluated by calculating the electron affinity of a large molecular system, Zn-protoporphyrin, which has previously been studied in the context of electron attachment. [17] The molecular geometry was taken from the Supplementary Material of Ref. 17. The calculation was performed on a dedicated workstation equipped with an Intel Xeon E5-2667 v4 processor (3.20 GHz) and 512 GB of RAM. The aug-cc-pVTZ basis set together with the aug-cc-pVTZ/C auxiliary basis was employed for the Zn atom, while the aug-cc-pVDZ basis set with the aug-cc-pVDZ/C auxiliary basis was used for all other atoms. Zn-protoporphyrin consists of 75 atoms and 326 electrons, with a total of 1184 virtual orbitals. Applying an SS-FNO truncation threshold of $10^{-4}$ reduced the virtual orbital space to 807 orbitals. Using an NAF threshold of $10^{-2}$ further reduced the number of auxiliary functions from 4053 to 3440. The total computation time for the SS-FNO-EA-ADC(3) calculation was 1 day, 10 hours, and 13 minutes, with the SCF step requiring only 18 minutes. The canonical ADC(2) calculation for a single root required 2 hours and 8 minutes, of which 34 minutes were spent on intermediate generation. In the truncated basis, the ADC(3) calculation took 1 day, 7 hours, and 20 minutes, with 17 hours and 49 minutes required for intermediate construction. The ADC(2) calculation took 20 minutes in the SS-FNO basis. The SS-FNO-EA-ADC(3) method yielded an electron affinity of 1.041 eV for the lowest electron-attached state. The corresponding Dyson orbital for Zn-protoporphyrin is depicted in Fig. 10.
@@ V Conclusion
We have developed, implemented, and benchmarked a reduced-cost ADC(3) approach for computing electron affinities based on the state-specific frozen natural orbital (SS-FNO) framework within the non-Dyson intermediate-state representation (ISR) formalism. Unlike the conventional MP2-based FNO scheme, the state-specific FNO approach selects an optimal subset of virtual natural orbitals tailored for each electron-attached state, thereby achieving a substantial reduction in computational cost for ADC(3) calculations. To further enhance efficiency, the density-fitting approximation was employed to eliminate the need for storing four-index electron-repulsion integrals. Additional truncation of the auxiliary basis using natural auxiliary functions (NAFs) provides further computational savings.
@@ Acknowledgements.
This work has been supported by the ANRF India((Project No. CRG/2023/002558). AKD acknowledges the research fellowship funded by the EU NextGenerationEU through the Recovery and Resilience Plan for Slovakia under project No. 09I03-03-V04-00117.